### v1

@@ Abstract
The evaluation of general domain, math-specific, and multimodal LLMs highlights the challenges presented by U-MATH. Our findings reveal that LLMs achieve a maximum accuracy of only 63% on text-based tasks, with even lower 45% on visual problems. The solution assessment proves challenging for LLMs, with the best LLM judge having an F1-score of 80% on -MATH
@@ 3 U-MATH
Accuracy is the primary performance metric for U-MATH, its text-only problems (U-MATH) and problems that include a visual component (U-MATH). The main performance measure for -MATH is macro-F1.
We use an LLM as a judge (Zheng et al., 2023) to measure the accuracy of the free-form answers against the golden solutions. A problem is considered solved only if all required questions are answered and all requested items (e.g., all saddle points) are correctly identified.
@@ 3.2 Dataset Statistics
The U-MATH benchmark comprises 1,100 carefully curated and validated mathematical problems. These problems are distributed across 6 core subjects with about 20% of the tasks incorporating visual elements, such as graphs, tables, and geometric figures, mirroring the multi-modal nature of real-world mathematical problems: Precalculus (Review), Algebra, Differential Calculus (+Differential Equations), Integral Calculus, Multivariable Calculus, and Sequences & Series.
@@ 4.2 U-MATH Results
Model U-MATH U-MATH Algebra Diff. C. Integral C. Multivar C. Precalculus Seq.& Series T V T V T V T V T V* T V* T V* 900 200 150 30 150 70 150 58 150 28 150 10 150 4 Text-only models Mathstral 7B 18.0 20.7 6.0 51.3 6.7 4.0 10.0 1.3 1.7 8.0 3.6 48.7 10.0 10.7 0.0 NuminaMath 7B 19.2 22.8 3.0 62.7 0.0 4.0 7.1 1.3 0.0 6.0 3.6 51.3 0.0 11.3 0.0 Llama-3.1 8B 22.3 26.1 5.0 59.3 3.3 6.7 5.7 9.3 3.4 11.3 3.6 54.7 10.0 15.3 25.0 Qwen2.5 7B 33.8 40.0 6.0 86.0 10.0 12.7 1.4 10.0 12.1 26.7 3.6 75.3 0.0 29.3 0.0 Qwen2.5-Math 7B 38.4 45.2 7.5 87.3 6.7 18.7 5.7 8.0 10.3 36.0 10.7 80.7 0.0 40.7 0.0 NuminaMath 72B 25.0 29.7 4.0 74.7 3.3 6.7 4.3 4.0 3.4 11.3 3.6 62.7 10.0 18.7 0.0 Llama-3.1 70B 28.5 33.7 5.0 82.0 3.3 10.7 5.7 4.0 5.2 14.0 3.6 64.0 0.0 27.3 25.0 Llama-3.1 Nemotron 70B 31.4 37.4 4.0 84.0 0.0 14.7 2.9 4.0 3.4 25.3 7.1 64.0 20.0 32.7 0.0 Qwen2.5 72B 41.0 48.6 7.0 88.7 6.7 22.7 4.3 12.0 6.9 40.0 17.9 83.3 0.0 44.7 0.0 Athene-V2 72B Chat 46.2 54.6 8.5 88.7 3.3 34.0 4.3 16.0 6.9 50.7 21.4 88.7 10.0 49.3 50.0 Qwen2.5-Math 72B 50.2 59.0 10.5 92.7 6.7 35.3 7.1 20.7 17.2 58.0 7.1 90.0 0.0 57.3 50.0 Multimodal models Pixtral 12B 15.5 15.6 15.5 44.7 23.3 1.3 34.3 0.7 0.0 3.3 0.0 32.0 0.0 11.3 0.0 Llama-3.2 11B Vision 17.0 18.6 10.0 54.0 10.0 1.3 20.0 1.3 1.7 4.7 3.6 43.3 10.0 6.7 0.0 LLaVA-OV Qwen2-7B 17.7 20.7 4.5 60.7 6.7 4.0 5.7 1.3 1.7 5.3 3.6 43.3 10.0 9.3 0.0 Qwen2-VL 7B 20.4 21.4 15.5 62.7 10.0 4.7 32.9 0.7 5.2 6.7 7.1 45.3 0.0 8.7 0.0 Qwen2-VL 72B 31.2 32.2 26.5 80.7 26.7 9.3 40.0 2.0 13.8 14.7 28.6 65.3 10.0 21.3 0.0 Llama-3.2 90B Vision 32.6 36.3 16.0 85.3 26.7 10.7 25.7 2.7 1.7 22.7 7.1 65.3 20.0 31.3 25.0 Claude Sonnet 3.5 35.1 36.1 30.5 76.0 33.3 12.0 41.4 7.3 17.2 21.3 28.6 65.3 30.0 34.7 25.0 GPT-4o-mini 37.2 40.3 23.0 88.0 16.7 16.7 31.4 4.0 10.3 24.0 35.7 77.3 20.0 32.0 25.0 GPT-4o 43.5 46.4 30.0 91.3 30.0 18.7 32.9 10.0 20.7 41.3 42.9 79.3 30.0 38.0 25.0 Gemini 1.5 Flash 51.3 53.8 40.0 91.3 50.0 36.0 45.7 14.0 24.1 44.0 50.0 80.7 30.0 56.7 50.0 Gemini 1.5 Pro 60.1 63.4 45.0 91.3 60.0 50.7 47.1 27.3 24.1 60.7 57.1 87.3 70.0 63.3 50.0
@@ Acknowledgement
We would like to give special thanks to the dedicated team of Gradarius experts who played a crucial role in validating problems and ensuring their quality: Jan Cannizzo Paul Schwartz Andrey Nikolaev Arina Voorhaar Chloe Weiers Funda Gul Tatiana Zenkevich Igor Teplukhovskiy Anastasiia Feklina Ruslan Akmetdinov Ekaterina Eremina Sofia Tekhazheva

### v3

@@ Abstract
The evaluation of general domain, math-specific, and multimodal LLMs highlights the challenges presented by U-MATH. Our findings reveal that LLMs achieve a maximum accuracy of only 63% on text-based tasks, with even lower 45% on visual problems. The solution assessment proves challenging for LLMs, with the best LLM judge having an F1-score of 80% on -MATH.
@@ 3 U-MATH
Accuracy is the primary performance metric for U-MATH, its text-only subset (U-MATH) and the subset of problems including a visual component (U-MATH). We use an LLM-as-a-judge (Zheng et al., 2023) to measure the accuracy of the free-form answers against the golden solutions. A problem is considered solved only if all required questions are answered and all requested items (e.g., all saddle points) are correctly identified.
@@ 3.2 Dataset Statistics
The U-MATH benchmark comprises 1,100 carefully curated and validated mathematical problems. These problems are distributed across 6 core subjects with about 20% of the tasks incorporating visual elements, such as graphs, tables, and geometric figures, mirroring the multi-modal nature of real-world mathematical problems: Precalculus, Algebra, Differential Calculus (+Differential Equations), Integral Calculus, Multivariable Calculus, and Sequences & Series.
@@ 4.2 U-MATH Results
Model U-MATH U-MATH Algebra Diff. C. Integral C. Multivar C. Precalculus Seq.& Series T V T V T V T V T V T V* T V* 900 225 150 30 150 80 150 68 150 28 150 13 150 6 Text-only models Mathstral-7B-v0.1 16.36 20.1 1.3 56.0 3.3 4.0 1.5 2.0 1.2 6.0 0.0 40.7 0.0 12.0 0.0 Llama-3.1-8B 15.47 19.0 1.3 52.0 3.3 4.7 2.9 1.3 0.0 10.0 0.0 36.7 0.0 9.3 0.0 Qwen2.5-7B 32.18 39.8 1.8 77.3 3.3 14.0 0.0 10.7 1.2 34.7 7.1 72.7 0.0 29.3 0.0 Qwen2-Math-7B 28.62 35.1 2.7 77.3 3.3 9.3 1.5 8.0 2.5 23.3 7.1 70.7 0.0 22.0 0.0 Qwen2.5-Math-7B 33.24 41.0 2.2 64.0 0.0 16.7 1.5 21.3 0.0 35.3 7.1 72.0 7.7 36.7 16.7 Llama-3.1-70B 26.13 32.0 2.7 70.0 10.0 10.0 0.0 5.3 2.5 25.3 3.6 55.3 0.0 26.0 0.0 Qwen2.5-72B 34.04 41.9 2.7 69.3 0.0 25.3 0.0 10.7 1.2 39.3 14.3 70.0 0.0 36.7 16.7 Qwen2-Math-72B 35.20 43.6 1.8 80.0 0.0 22.7 0.0 16.0 2.5 32.0 0.0 76.7 0.0 34.0 33.3 Qwen2.5-Math-72B 41.16 50.6 3.6 73.3 3.3 33.3 0.0 23.3 7.5 48.7 3.6 82.0 0.0 42.7 0.0 Multimodal models Qwen2-VL-7B 17.33 20.3 5.3 51.3 6.7 8.7 10.3 1.3 1.2 6.0 3.6 44.0 7.7 10.7 0.0 LLaVA-OV(Qwen2-7B) 14.40 17.8 0.9 48.7 0.0 4.7 1.5 0.7 0.0 7.3 3.6 36.7 0.0 8.7 0.0 Pixtral-12B-2409 15.64 18.1 5.8 46.7 16.7 4.7 10.3 0.7 1.2 6.7 0.0 41.3 0.0 8.7 0.0 Qwen2-VL-72B 26.93 30.1 14.2 70.0 13.3 11.3 20.6 5.3 7.5 18.7 21.4 58.7 15.4 16.7 0.0 GPT-4o 36.53 41.8 15.6 80.0 16.7 22.0 14.7 11.3 13.8 38.7 25.0 68.0 15.4 30.7 0.0 Gemini-1.5-Pro 48.89 53.4 30.7 84.7 56.7 37.3 27.9 27.3 22.5 42.7 28.6 81.3 46.2 47.3 16.7
@@ Acknowledgement
Some of the problems in U-MATH and mu-MATH are sourced from OpenStax under CC BY 4.0.
@@
We would like to give special thanks to the dedicated team of experts from Gradarius and Stevens Institute who played a crucial role in validating problems and ensuring their quality: Jan Cannizzo Paul Schwartz Andrey Nikolaev Arina Voorhaar Chloe Weiers Funda Gul Tatiana Zenkevich Igor Teplukhovskiy Anastasiia Feklina Ruslan Akmetdinov Ekaterina Eremina Sofia Tekhazheva